OpenStudy (anonymous):

The time required to finish a test is normally distributed with a mean of 60 minutes and a standard deviation of 10 minutes. What is the probability that a student will finish the test between 40 and 70 minutes? Some one help!

OpenStudy (anonymous):

Z score = ( 40 - 60 )/ 10 = - 2 Z score = ( 70 - 60) / 10 = 1 P ( 40 < X < 70 ) = P ( -1 < Z < 1 ) = ... Look up your Z score table to get Probability

OpenStudy (anonymous):

@Stina_rox96 Can you continue?

OpenStudy (kropot72):

The limits are -2 sigma and +1 sigma from the mean. The probability of being -2 sigma from the mean is 0.4773 and the probability of being +1 sigma from the mean is 0.3413. What is the overall probability?

OpenStudy (kropot72):

@Chlorophyll Please check your final equation.

OpenStudy (anonymous):

I mean = P ( -2 < Z < 1 ) = ...

OpenStudy (kropot72):

@Stina_rox96 You will get the same answer with either method. Mine relies on known properties of the Standard Normal Distribution Curve